OpenStudy (anonymous):

2x-y-6=0 find the x and y-intercepts

OpenStudy (anonymous):

what do you know?

OpenStudy (anonymous):

so far this is what I've done y+2x-6=0

OpenStudy (anonymous):

idk how to start

OpenStudy (anonymous):

when you solve for x and y intercepts you substitute the other variable with 0 so if you were to solve for x you would replace y with 0 and if you were to solve for y you would replace x with 0

OpenStudy (zzr0ck3r):

when x is 0 we have -y-6 = 0 or y = -6 so the y inttercept at (0,-6) when y = 0 we have 2x-6 = 0or x = 3 so the x intercept is (3,0)

OpenStudy (anonymous):

ok thats cool. lets see. first what would be the best way the formula should be in graph . ??

OpenStudy (zzr0ck3r):

note: y intercept is when x is zero and x intercept is when y is 0

OpenStudy (anonymous):

yea, ok. thank you everyone. I apreciate it

OpenStudy (anonymous):

guys concept first before solving. !!!

OpenStudy (anonymous):

a little too late for that... I always work the problem out myself anyways though

OpenStudy (anonymous):

look think of what are you trying to get first and then go about learning the mechanical side of it.

OpenStudy (anonymous):

good luck.

OpenStudy (anonymous):

yea

OpenStudy (zzr0ck3r):

I like to se it done, and then I can get the concept.... its fine.

OpenStudy (anonymous):

exactly what I'm doing...

OpenStudy (anonymous):

so where is the concept ?

OpenStudy (zzr0ck3r):

read up. no need to repeat

OpenStudy (zzr0ck3r):

note: y intercept is when x is zero and x intercept is when y is 0

OpenStudy (anonymous):

everybody doesn't learn the same ways so relax

OpenStudy (anonymous):

no nothing up ??

OpenStudy (zzr0ck3r):

its fine, if peole dont want to learn(not saying you dont) we cant force them to. I understnad what your saying, and if they ask the same sort of question over and over then Im with you. but this is not the case so far here.

OpenStudy (zzr0ck3r):

note: y intercept is when x is zero and x intercept is when y is 0 ^^^^^^^^^^^ this was all up...

OpenStudy (zzr0ck3r):

its ok man

OpenStudy (anonymous):

i actually worked the problem out instead of just writing what he said so I can understand how he got his answer

OpenStudy (anonymous):

thank you. i think people will be better off if they truly understand whats happing behind the equation than simply to solve it .

OpenStudy (zzr0ck3r):

again she is saying it helped her.

OpenStudy (anonymous):

my answer obviously helped the most :D